Session: Sep 2000 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Huang Xinyong at Fri Sep 15 12:42:27 METDST 2000) Success

Torun (K.Borkowski at Fri Sep 15 18:56:26 METDST 2000) Minor failures
During the experiment at times FS reported "no_1pps" in BBC1 or BBC2. Also, probably before the expt. started there occurred a 1s jump in the formatter time, which has been corrected at about 13:30 by addition of 1 s to the formatter time.

Effelsberg (A. Kraus at Sat Sep 16 18:09:57 METDST 2000) Minor failures
Few times late on source.

Westerbork (Tony Foley at Mon Sep 18 09:38:19 METDST 2000) Minor failures
Several scans outside hour angle range replaced by local calibrators. Also lost from 07:26 to 07:32 because of a local scheduling error Used tied array with 14 telescopes

Onsala (A. Polatidis at Mon Sep 18 23:42:26 METDST 2000) Success
No obvioys problems

Urumqi (Liu Xiang at Wed Sep 20 10:32:36 METDST 2000) Success

Noto (C.Trigilio at Wed Sep 20 14:09:58 METDST 2000) Success

Jodrell_Bank (Alastair Gunn at Tue Sep 26 06:03:20 METDST 2000) Success
FS restart at about 12:14 UT, LOs reset but apparently no loss of data. Another restart at 12:29 UT. Data should be fine. No other known problems.